OpenStudy (anonymous):

If g(4) = 16, g(5) = 18, g(6) = 20, then g(3) = _____?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

the common ratio between the terms is 2, therefore, 16-2= 14. .

OpenStudy (anonymous):

simplest way to explain this how did u get 14?

OpenStudy (anonymous):

medal me!

OpenStudy (anonymous):

each equation increases by 2 so logically the next number should be 22

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Yes, g(3) = 14

OpenStudy (anonymous):

g(x)=2x+8 g(3)=2(3)+8=14

OpenStudy (anonymous):

it is 14

OpenStudy (anonymous):

before g(4) is 14 after g(7)=22

OpenStudy (anonymous):